### v1

# Robotic deployment on construction sites: considerations for safety and productivity impact

Abstract
Deploying mobile robots in construction sites to collaborate with workers or perform automated tasks such as surveillance and inspections carries the potential to greatly increase productivity, reduce human errors, and save costs. However ensuring human safety is a major concern, and the rough and dynamic construction environments pose multiple challenges for robot deployment. In this paper, we present the insights we obtained from our collaborations with construction companies in Canada and discuss our experiences deploying a semi-autonomous mobile robot in real construction scenarios.
I INTRODUCTION
The imperative of ensuring human safety in shared workplaces with robotic systems presents a major challenge for the field deployment of such systems. The risks associated with human-robot collaboration encompass both direct and indirect hazards, with direct risks arising from collisions between the platform and obstacles such as equipment or personnel, while indirect risks stem from environmental factors affecting robot functionality, including noise and wind, as well as potential distractions posed by the presence of the robot itself. Addressing these concerns, Moud, Hashem Izadi, et al. proposed a model for evaluating the safety of mobile robot deployment based on physical characteristics and environmental considerations [1].
In a similar vein, Augustsson et al. advocated for the establishment of defined safety zones surrounding robots to facilitate safe interaction with humans, with the flexibility of zone geometries tailored to accommodate varying work scenarios [2]. Complementing this approach, Truong et al. emphasized the importance of human safety in shared environments, implementing a system wherein robots dynamically generate safe navigation zones upon detecting individuals through sensor feedback [3]. Furthermore, Shin et al. proposed a framework aligned with ISO/TS 15066 to evaluate the collision peak pressure of collaborative robots in proximity to humans, underlining the criticality of safety considerations within shared workspaces [4].
Our work focuses on the construction industry, recognized for its inherent hazards to human workers. It presents a particularly challenging environment for the deployment of autonomous mobile robots. Kim et al. highlighted persistent security concerns surrounding human-robot collisions in construction sites, accentuated by the mobility of robots [5]. While early endeavors in human-robot collaboration within construction settings have demonstrated potential efficiency gains, the complexities of maintaining safety alongside productivity remain significant [6].
The objective of this paper is to study the safety and productivity impacts of the deployment of autonomous mobile robots on construction sites, as perceived in a real industry context. For this purpose, we first review the industry standards on this subject, followed by directed interviews with construction industry experts in mobile robotics. Finally, we propose a mobile robot prototype specifically equipped for construction site progress monitoring.
Figure 1: Our autonomous mobile robot, Journeybot, navigating a construction site
II Industry Standards
Among the most recent standards, ANSI/RIA R15.08 stands as a pivotal framework for regulating industrial mobile robots (IMRs). Within its purview, both autonomous mobile robots (AMRs) and Automated Guided Vehicles (AGVs) are encompassed under the IMR designation when deployed within industrial settings. However, AGVs and AMRs exhibit distinct operational scopes: while AGVs adhere to predetermined routes, AMRs leverage sensor technologies to navigate around obstacles autonomously. Rooted in foundational principles from ANSI/RIA R15.06 and ANSI/ITSDF B56.5, which respectively address industrial robot safety and guided industrial vehicles, ANSI/RIA R15.08 stipulates comprehensive requirements for IMRs and their associated attachments. Noteworthy provisions include specifications for operational modes, control functionalities, and presence-sensing devices, among others.
In tandem with adherence to industry standards, due consideration must also be accorded to local regulations pertinent to the targeted workplace environment. National and local codes and regulations supersede industry standards, mandating strict compliance, particularly in construction sites where Health and Safety Standards hold paramount importance. Each construction site is overseen by a Safety Agent tasked with ensuring adherence to these regulations.
On a construction site, the traffic plan and traffic control are essential, as they constitute a measure to prevent unexpected contact between pedestrians and vehicles. To this end, the project manager must plan the movement of the latter to restrict backing maneuvers and put in place safety measures to protect any person circulating on the site. They must also inform, beforehand, any person who has to circulate on the site of the safety measures provided for in the circulation plan (CSTC, art. 2.8 and 2.8.2). This traffic plan must obviously include any IMR likely to be deployed. The site signaller must meet certain conditions, in particular: wear high-visibility fluorescent yellow-green safety clothing, remain visible to vehicle drivers, and stay out of the latter’s path (CSTC, art. 2.8.4). Therefore, any worker must remain visible and identifiable to an IMR by wearing high visibility clothing.
Self-propelled vehicles must be equipped with effective brakes and warning devices to be used when approaching pedestrians, doors, turns, and dangerous locations. This requirement does not apply to mechanical rams mounted on crawlers, skidders, and all-terrain vehicles (CSTC, art. 3.10.2). All equipment must only be used by an experienced driver or under their supervision (CSTC, art. 3.10.4). Finally, all self-propelled vehicles must be equipped with an audible warning device which must be installed within reach of the driver, have a sound specific to the type of equipment and different from any other signal from the same construction site and have sufficient noise level to dominate construction noise (CSTC, art. 3.10.12).
III Insights from the industry
Mobile robots have a wide range of usages for construction work, with drones (mainly outdoor ones) by far the most popular robots currently deployed. Several of their applications outdoor are envisioned for future indoor mobile robots, such as mapping, surveying, inspections, and marketing purposes. Indoor ground robots would be mainly used for progress monitoring and quality control. Other more specialized robots - equipped with IoT receivers, or able to draft on a concrete slab - are tested in various R&D projects.
Thanks to our close collaboration with two major companies in Canada already deploying robotic systems on their construction sites, we tackle some of the main challenges through directed interviews.
III-A Accessibility and acceptance
Robots are deployed most commonly on a bi-weekly or monthly basis on most of the construction sites, and again, mostly flying robots.
The adoption of robotic systems by workers and the general public varies according to the type of robot used. Drones are becoming more commonly accepted, quite often already seen as common industrial equipment. However, ground robots that operate independently alongside humans are still noticed and trigger skepticism. Workers often criticize ground robots and are afraid of being displaced by them. This significantly affects their acceptance on building sites. The appearance of the ground robots plays a significant role in their acceptance by human workers. More industrial designs are less criticized and better welcomed by the worker community than life-like ones (dog, spider, humanoid, etc).
III-B Deployment Logistics, Operations & Impact on Schedule
Regulatory requirements, technological limitations, and location-specific variables all have an impact on logistical issues in the deployment of robotic systems. Drones have become a vital part of the construction industry and are no longer considered as an external or exotic element on construction sites, thus having less deployment complexity. However, the same cannot be said for the ground robots as they are still seen as external elements on a construction site and pose challenges with regulations and acceptance by workers. When considering the autonomy of operations in industry, a notable distinction emerges between ground robots and drones. Ground robots, more often than not, rely on teleoperation, where human operators guide their movements and actions remotely. This is mandated by the regulations when there are human workers on site. Conversely, drones predominantly operate autonomously with waypoint navigation. This autonomy allows drones to execute predefined tasks efficiently, such as surveillance, mapping, or delivery, with minimal human intervention, however, human supervision is mandatory.
During the adoption and developmental phase of robots in construction, the deployment on construction sites may have significant impacts in terms of logistics and impacts on schedules. However, a successful integration of robotics on construction sites can result in increased productivity and safety.
III-C Suitability of the construction sites for robotic deployment
Construction sites differ in their viability for robotic deployment, with building projects being more suitable for drone use due to outside conditions and fewer complicated impediments. Indoor projects involve problems such as increased complexity, greater interaction with people, and security issues, making them less suitable for robotic deployments. However, developments in technology and legal frameworks may provide potential for indoor robotics in the future.
Successful robotic system deployments are dependent on a variety of elements, including technology developments, compliance with safety laws, public trust, and good communication. Advances in battery life and navigation technologies will improve task efficiency and worker protection. Compliance with safety rules guarantees the safe operation of drones and robots, which will promote acceptability among site people. Building public trust via openness and resolving concerns is critical to the widespread deployment of robotic technology in the building sector.
IV Autonomous mobile robot prototype deployment
Based on our long-standing collaboration with construction companies already owning robotic systems and testing their deployment on construction sites, we developed a prototype robot, Journeybot, to assess the complexity of these field deployments.
IV-A The Robot
We developed an autonomous mobile robot prototype to perform deployments in construction sites and collect sensor data, named Journeybot. Our robot is built from a four-wheeled unmanned ground vehicle, the Clearpath Jackal, equipped with a hybrid vision/laser scan sensing system as shown in Fig. 2. This robotic system is capable of collecting data while the robot is moving and can navigate semi-autonomously, with a remote operator sending high-level commands.
Figure 2: Mobile robot platform equipped with various sensors
The sensing suite of the Journeybot is designed to facilitate comprehensive data collection for digital twin applications. Strategically positioned LiDARs (2) and depth cameras (5) grant a panoramic coverage of the robot’s surroundings for robust environmental perception.
The robot’s pose in the map is obtained through the use of a ROS implementation of the Adaptive Monte Carlo localization algorithm from [7]. Before deploying the robot, wall geometry information is extracted from the Building Information Model (BIM) to generate an occupancy grid of the building. During the robot navigation, this map, the odometry, and the laser scan data from the horizontally mounted Velodyne LiDAR are fed to the localization algorithm, which then estimates the robot’s current pose in that map. When a desired pose command is sent by the remote operator to the robot, a navigation algorithm calculates the shortest path to that pose that avoids all obstacles known from the occupancy grid. This path is constantly updated as the robot moves and new obstacles are detected. More details are given in [8].
IV-B The Fields
Pomerleau Construction based in Montreal, Canada, opened three construction sites to our team, shown in Fig. 3. Each deployment presented unique environmental challenges:
  • Hospital Building, Office Floor: In the later stages of construction, this floor consisted of long, narrow corridors almost clear of construction equipment and materials, although numerous workers were moving around, some occasionally crossing paths with the robot.
  • Hospital Building, Mechanical Floor: In stark contrast to the office floor, this deployment unfolded on a floor dedicated to heating and ventilation equipment. The space was filled by large boilers, electrical generators, water pumps, and piping systems. Despite the absence of construction workers, narrow passages between equipment were challenging.
  • Second Building, Early Construction Stage: This early-stage construction site was littered with construction materials and equipment. We deployed the robot on the second floor, where worker presence was minimal, in a vast yet cluttered navigation space. Piles of construction material and assorted obstacles demanded agile maneuvering of the robot.
  • Third Building, Middle stage: This factory plant was in the middle stages of construction, with partially installed production lines and scattered construction materials. The larger spaces offered ample room for the robot to navigate easily.
Figure 3: Journeybot in different construction sites: (a) Hospital Building, Office Floor; (b) Hospital Building, Mechanical floor; (c) Second Building, Early Construction Stage
Each of these deployments was preceded by a careful analysis by the team of the potential challenges on site.
IV-C The Logistics
IV-C1 Preparation
Several preparation steps preceding each deployment were instrumental in deploying safely our robotic platform into the dynamic construction environments:
  • Deployment plan validation: On the days preceding each deployment, we conducted comprehensive testing of the robot’s navigation and data collection capabilities within the controlled environment of our university laboratory. We tailored each test from the field deployment layout - trying to reproduce similar trajectories and environments, as much as possible. This involved deploying the robot in several different locations on the campus. Subsequently, we ensured that both the robot’s battery and the battery of the remote controller were fully charged and ready for deployment.
  • Collision Avoidance System testing: Construction site environments require safe management of both equipment and personnel in the robot vicinity. Our autonomous navigation stack includes an emergency collision avoidance system, halting the robot upon detecting anything too close. Rigorous tests are conducted before each deployment to confirm its efficacy.
  • Assessment of Terrain Compatibility: Our team carefully studied the construction site topography including slopes, stairs, openings, etc. We conducted extensive simulations to confirm the robot’s capability to traverse such terrains. From these simulations, we devised contingency plans to manually navigate the robot over obstacles exceeding its capacity.
These preparatory measures not only optimized the functionality of our technology but also underscored our commitment to safety and efficiency in real-world applications. These steps may not provide direct insights into research novelty but they help the research team better assess the reality of the application of their research and they ensure better acceptance of the technology on the field (less prone to failure).
IV-C2 On site
Upon arriving at the construction sites, we first performed the following procedure:
  • Present Journeybot to the various stakeholders (workers, supervisors) on the site for them to get to know the new mobile system that will now share their workspace. This step is critical to make sure they know the risks and how to behave in some specific situations with this AGV to reduce the probability of incidents. We let some employees operate the robot and see the sensors’ output on its management interface.
  • Walk through the construction site to survey and validate the path plan and topology studied. On complex fields, we also operate the robot manually for the first run before launching the autonomous navigation stack. It is essential to confirm the functionality of the data acquisition system and that Journeybot is adapted to this environment.
During the actual deployment, we sent commands to the robot to explore the environment autonomously and monitor it at all times. One team member walked alongside Journeybot while monitoring its sensors and navigation software output from a laptop that was being carried by one of the engineers. Another team member held the remote controller to take over the robot control if necessary. We would also constantly monitor the environment itself: look out for obstacles and construction workers moving through the environment.
V CONCLUSIONS
Ensuring human safety in shared workplaces with robotic systems presents This paper presented our endeavors to study safety considerations and productivity impacts of deploying autonomous mobile robots in construction sites. Thanks to our collaborations with major construction companies in Canada we obtained insights on the subject from the industry perspective and performed deployments in real scenarios, allowing us to identify key points and strategies that we believe are very helpful in these situations. The main key points are: performing a thorough assessment of the site before the deployment, to identify possible challenges for the robot; testing the robotic system in simulations and physical environments that recreate the main characteristics and challenges of the real construction site; maintaining checklists to be verified after each deployment to avoid mistakes and increase consistency. In the future, these protocols can be applied with other robotic platforms and extended for more varied environments.
ACKNOWLEDGMENT
The author thanks CANAM and Pomerleau Construction employees for their precious insights and their generosity in sharing their experience with field deployment of mobile robotic systems.